46. Using Hesss law, calculate the AH value for the following reaction: (8 marks) FeOce) + CO(g) → Fe(s) + CO2(g) Use theseI need clear solutions and step by step answer. would be helpful to explain what you did too.
Answer #1

Please carefully follow the answer below.

46.

Feot co Fe co2 Rte t 3 C02 A 25-0kJ 3 Peao t o , 4M40 KT 304Co 3 Feot Uoz 0y t o (3) 44 58.0 us He las, Bui thi three euahin,the whou equahan b 6, we get Now 6x t Co2 kJ 38 An Gteot 3Fea 2 +7co 2tegOu +b Fe tWz step 3 Now Sub tout the eguahon fom thBalcmeod thermo cheswiča equaten 47 a Natto3 ) Na Co3 Lt Co, g)+0 ) Frast all A for Na,co ) = -1130-7 k)/ma = 393-5 KJ/mol A